The demand function of goods is Q=100-P+2Y+A ,where Q,P,Y and A donate Quantity demand ,price income and advertisement expenditure respective . calculate the demand when P=40,Y = 40 and A = 6.assume that price and income are fixed , calculate additional expenditure needed to raise demand to 179


1
Expert's answer
2021-03-10T19:36:12-0500

(a)

"Q=100-40+2\cdot40+6=146."

(b)

"A=Q-100+P-2Y,""A=179-100+40-2\cdot40=39."